Solve x2 - 2x - 24 = 0 by using completing the square.
Answer by Boreal(15235)

x^2-2x=24, moving the constant to the other side.
Take half of the 2 and square it.
x^2-2x+1=24+1, adding 1 to both sides
(x-1)^2=25 or 5^2
take the square root of both sides
x-1=+/-5
x=1+/-5; the roots are 6 and -4
36-12=24
16+8=24